3. (4 pts) Using a normal distribution having mean of 200 and standard deviation of 15.0, calculate the following probabilities:

a. P(X <= 180)

b. P(X<=220)

c. P(X >= 220)

d. P(180 <= X <= 220)
Answer #1

3

mean (\mu) = 200

standard deviation (\sigma) = 15
a)

\\ P(x\le180)\\=P(z \le \frac{x-\mu}{\sigma})\\=P(z \le \frac{180-200}{15})\\=P(z\le-20/15)\\=\mathbf{0.0912}

b)
\\ P(x\le220)\\=P(z \le \frac{x-\mu}{\sigma})\\=P(z \le \frac{220-200}{15})\\=P(z\le20/15)\\=\mathbf{0.9088}

c)
\\ P(x\ge220)\\=P(z \ge \frac{x-\mu}{\sigma})\\=P(z \ge \frac{220-200}{15})\\=P(z\ge20/15)\\=\mathbf{0.0913}

d)
\\ P(180 \le x\le220)\\\\=P(\frac{x-\mu}{\sigma} \le z \le \frac{x-\mu}{\sigma})\\\\=P(\frac{180-200}{15} \le z \le \frac{220-200}{15})\\\\=P(-20/15 \le z\le20/15)\\\\=\mathbf{0.8175}

please upvote